A student has a mass of 49.2 kg and is sitting on a seat of a rollercoaster. The ride makes the seat move up and down. Determine the normal force acting on the student if they are: a) moving at a constant velocity of 10.0 m/s b) accelerating 2.1 m/s ^ 2 (down) 


1
Expert's answer
2021-01-07T10:55:30-0500

a)


"N=mg=(49.2)(9.8)=482\ N"

b)


"N=m(g-a)=(49.2)(9.8-2.1)=379\ N"